Lincoln Financial is considering an investment proposal that will require the purchase of $80,000 of machinery and initial investment in working capital of $40,000. The proposal is expected to yield cash flows and income (before depreciation and taxes) of $48,000 annually for four years. Depreciation is to be taken for tax purposes using the straight-line method and no salvage value. The investment in working capital will be recovered at the end of four years. The company will not accept a proposal that will not earn at least 15%. The present value of an annuity of $1 for four periods at 15% is 2.85498; the present value of $1 due in four periods at 15% is 0.57175. The tax rate in effect is 40%. Determine whether the project is acceptable using the net present value method.
Answer #1

NPV=-80000-40000+40000/1.15^4+((48000-80000/4)*(1-40%)+80000/4)/15%*(1-1/1.15^4)
=-80000-40000+40000*(P/F,15%,4)+((48000-80000/4)*(1-40%)+80000/4)*(P/A,15%,4)
=-80000-40000+40000*0.57175+((48000-80000/4)*(1-40%)+80000/4)*2.85498
=7933.26400

Acceptable